Using the given reaction, answer the given questions: 2H2(g) + O2(g) → 2H2O(l) ∆H = -572 KJ

a. How much heat is liberated for the production of 1.00 mol of H2O(l)?

b. Determine the heat evolved when 4.13 g of hydrogen is reacted with excess oxygen.

c. Determine the amount of heat liberated when 190 g of oxygen is reacted with excess hydrogen. 


1
Expert's answer
2021-11-17T04:37:01-0500

A) H= -572KJ this is for 2moles so for one moles- 286KJ


B) 2H2 +O2 = 2H2O

  4g + 32g =36g SO delta H will be -572KJ


C) H2 +1/2O2 = H20

V = 2.00 × 10⁸ L

p = 1.00 atm

T = 25 °C

ΔH = -286 kJ·mol⁻¹

Calculations:

(a) Convert temperature to kelvins

T = (25.1 + 273.15) K = 298.25 K


(b) Moles of hydrogen

Use the Ideal Gas Law:

pV = nRT

n = (pV)/(RT)

n = (1.00 × 2.00 × 10⁸)/(0.082 06 × 298.25) = 8.172 × 10⁶ mol

(c) Heat evolved

q = nΔH = 8.172 × 10⁶ × (-286) = -2.34 × 10⁹ kJ

The hydrogen in the Hindenburg released 2.34e9 kJ .